A uniform magnetic field of induction B is changing in magnitude at a constant rate dB/dt.  copper wire of mass M is to be drawn into  wire of rdius r and formed into a circular loop of radius R. If P(rho) is the specific resistance and  delta( ) is the density of wire, then what is the induced current in the loop? Assume B is perpendicular to the loop..

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

Dear Saurabh, 

Resistance s=ρl/A,l=2πr, A=πr2, s=2ρR/r2, V= d/dt(πR2B)=πR2dB/dt, current I=V/s =(πRr2/2ρ)(dB/dt)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar.. 

Askiitians Expert

Chetan Mandayam Nayakar – IIT Madras